OpenStudy (anonymous):

how do i work a fraction 2/3 of 12 ?

OpenStudy (anonymous):

2*12 /3... 24/3=8

OpenStudy (anonymous):

is that wat u mean?

OpenStudy (anonymous):

OF always means multiply so it is 2/3 times 12 then calculate it

OpenStudy (anonymous):

of means multiply, so you have to multiply 2/3 as a decimal or do it as a fraction, so 2/3*12/1=24/4. and when you simplify that, it's 8. so 2/3 of 12 is 8.